# Spin-1/2 kagome Heisenberg antiferromagnet: Machine learning discovery of the spinon pair density wave ground state

### v1

@@ II Wave-function representation with group convolutional neural networks
and $\vec{\sigma}=\left\{\sigma(\vec{r}_{1}),...,\sigma(\vec{r}_{N})\right\}$ with $\vec{r}_{1},...,\vec{r}_{N}$ being positions of the lattice sites. In other words, spin configurations $|\vec{\sigma}\rangle$ are the network inputs, and the complex coefficients $\psi(\vec{\sigma})$ are the network outputs as illustrated in FIG. 2.
@@
All $N_{l}$ hidden layers are followed by the $\Gamma$ activation function. To obtain $\psi(\sigma)$ in equation (2) the final output layer exponentiates complex-valued feature maps composed of two real-valued feature maps, calculates sum of these exponentiated values, and projects the sum on a particular irreducible representation of the symmetry group $G$ that corresponds to the symmetry properties of the ansatz wavefuntion:
@@
Within group theory representations are eigenfunctions of the symmetry group operations. Irreducible representations are eigenfunctions of all group operations whilst reducible representations are eigenfunctions only of a subset of these operations. Each irreducible representation has different set of eigenvalues. These eigenvalues are characters $\chi_{g}$ in the equation (9). For non-degenerate symmetry groups where every irreducible representation consists of one eigenfunction, characters are $+1$ or $-1$, whilst for degenerate symmetry groups characters can differ from ±$\pm 1$.
@@
We also note that using local embedding kernels $K$, obtained by restricting nonzero $K$ elements only to local clusters, stabilizes the optimization procedure and allows numerical calculations with deeper neural networks.Roth ; Roth2 The optimization procedure, i.e. the network training where the network learns from the feedback from variational principle, has faster convergence for GCNN architectures with local kernels since the kernel locality structures the representation of long-range correlations and simplifies the learning landscape. Locality in the convolutional kernels $W$ is imposed by restricting the kernels to the symmetry operations of the form $\hat{g}=\hat{t}\hat{p}$ where $\hat{t}$ is a lattice translation and $\hat{p}$ is a lattice point-group symmetry operation.
In our calculations for the kagome antiferromagnet kernels $K$ are restricted to the symmetric local cluster illustrated in FIG. 3. We find that choosing larger clusters with different symmetric shapes does not improve convergence nor stability of the algorithm. Set of the crystal momenta related to the lattice translations and the lattice Brillouin zone are illustrated in FIG. 4. Positions of the kagome lattice sites are $\vec{R}_{L_{1},L_{2},k}=L_{1}\cdot\vec{a}_{1}+L_{2}\cdot\vec{a}_{2}+\vec{r}_{k}$ where $\vec{a}_{1}=\left(1,0\right)$ and $\vec{a}_{2}=\left(1/2,\sqrt{3}/2\right)$ are two primitive vectors of the lattice unit cell containing three inequivalent sites, and $\vec{r}_{k}$ describe positions of the sites within the unit cell: $\vec{r}_{1}=\left(1/2,0\right)$, $\vec{r}_{2}=\left(1/4,\sqrt{3}/4\right)$ and $\vec{r}_{3}=\left(3/4,\sqrt{3}/4\right)$. Two primitive basis vectors $\vec{b}_{1}=2\pi\left(1,-1/\sqrt{3}\right)$ and $\vec{b}_{2}=4\pi\left(0,1/\sqrt{3}\right)$ in the reciprocal lattice are calculated as $\vec{b}_{1}=2\pi\frac{\hat{O}_{R}\vec{a}_{2}}{\vec{a}_{1}\cdot\hat{O}_{R}\vec{a}_{2}}$ and $\vec{b}_{2}=2\pi\frac{\hat{O}_{R}\vec{a}_{1}}{\vec{a}_{2}\cdot\hat{O}_{R}\vec{a}_{1}}$ where $\hat{O}_{R}$ represents $\pi/2$ (anti-clockwise) rotation matrix.
Described procedure of constructing the wave-function ansätze equivariant with respect to a discrete symmetry group cannot straightforwardly be generalized to continuous symmetry groups like $SU(2)$ spin rotation symmetry. Finding efficient ways to introduce continuous symmetries is still an open research problem with one of the possible solutions being sampling in the basis of irreducible representations instead of spins.Vieijra ; Vieijra2 ; Misawa
III Training scheme and wave-function optimization
Within the VMC approach variational wavefunctions described by GCNN ansätze are optimized using stochastic reconfiguration method introduced by Sorella et al..Sorella1 ; Sorella2 ; Sorella3 ; Becca The goal of VMC is to find the optimal set of (complex) parameters $\vec{\alpha}=\left\{\alpha_{1},...,\alpha_{N_{p}}\right\}$ in an ansatz wavefunction of the form (2) that minimizes a suitably chosen loss function. The standard optimization technique is the gradient descent (GD) method within which the parameters are updated according to the gradients of the chosen loss function with the iterative update rule
@@ III Training scheme and wave-function optimization
where $t$ denotes iteration, $\eta$ is the learning rate (the step size) and $\mathcal{L}$ is the loss function. All numerical methods that attempt to find a global minimum using local information for iterative parameter updates generally exhibit problems related to the local minima in the loss function landscape and slow convergence. Various stochastic methods are therefore introduced to help mentioned numerical algorithms to escape local minima and improve convergence, for example AdamKingma and natural gradient descent (NGD) method.Amari ; Park ; Martens ; Dong ; Park2 ; Stokes2
In the equation (13) -$\vec{\nabla}_{\vec{\alpha}}\mathcal{L}$ represent directions in which the parameters can be updated to obtain the largest decrease of the loss function. The parameters are adjusted by a small step length, i.e. moved for a small distance in the parameter space, in the direction determined by the gradient to optimize the loss function. The standard GD and NGD methods differ by how a small distance is defined. Whilst in the standard GD distance is Euclidean distance that corresponds to the flat metric on the parameter manifold, NGD replaces the flat metric by a generally non-flat metric. NGD method takes into account that changes of all parameters can not be treated equally and that it is not always correct to define distance in terms of how much the parameters are adjusted. Instead a change for each parameter has to be scaled according to the effect of the change on the neural network output distribution.
@@
where $\Delta\tau$ is a small imaginary time step that corresponds to the learning rate $\eta$ in the equation (14). After each imaginary time step $\Delta\tau$ the wavefunction $|\psi^{\prime}\left(\vec{\alpha}\left(\tau\right)\right)\rangle$ is projected to the variational subspace by minimizing $d_{FS}^{2}$ with $|\phi\rangle=|\psi\left(\vec{\alpha}\left(\tau\right)\right)\rangle+\Delta\tau\partial_{\tau}\psi\left(\vec{\alpha}\left(\tau\right)\right)$ and $|\phi^{\prime}\rangle=|\psi^{\prime}\left(\vec{\alpha}\left(\tau\right)\right)\rangle$ in the equation (15) to obtain a new set of parameters $\vec{\alpha}_{t+1}\equiv\vec{\alpha}\left(\tau+\Delta\tau\right)$:
@@
The most often choice of the loss function in VMC is the variational energy:
@@
$H=J\sum_{\langle i,j\rangle}\vec{S}_{i}\cdot\vec{S}_{j},$ (20)
$G_{ii}\rightarrow G_{ii}+\epsilon_{1}G_{ii}+\epsilon_{2}$ (21)
provides suitable regularization at the later stages of the calculations when the variational energy is relatively close to the ground state energy. However the mentioned regularization scheme is often unstable at the beginning of the network training exhibiting run-away-like instabilities typical for non-stoquastic quantum Hamiltonians with rugged variational manifold landscape.Bukov For GCNN architectures similar behavior was previously noticed in studies of the frustrated $J_{1}$-$J_{2}$ model.Roth ; Roth2
@@
and $\psi\left(\vec{\sigma},\vec{\alpha}\right)$ is an unnormalized neural-network ansatz wavefunction. Within this approach the effective temperature is set to some initial $T=T_{0}$ at the start of the training and then gradually lowered to zero as the training proceeds according to a schedule function $T_{n}$ for the effective temperature in the $n$-th training step.
Similarly as for the $J_{1}$-$J_{2}$ model,Roth ; Roth2 in our kagome lattice antiferromagnet study the function $T_{n}$ of the form $T_{n}=T_{0}e^{-\lambda n}$ with $T_{0}=0.5$ and $\lambda=0.02$ the most effectively solves the problem with initial instabilities for a range of random initial parameter values. We however note that for larger lattice sizes we still find residual instabilities for some initial random parameter values that are prone to the instabilities and lead to unstable trajectories in the network parameters manifold. This dependence on the initial parameter values can in general appear for non-stoquastic quantum Hamiltonians in rugged neural network landscapes.Bukov The sum in Eq.(22) is over all basis states $\vec{\sigma}$ and at large values of $T$ at the beginning of the training the network tries to first learn states with similar amplitudes for all basis states. This in addition to the training stabilization also significantly increases achievable results accuracy which is consistent with findings in several previous neural network studies that indicate improved algorithm performance with including pre-training of the phases.Szabo ; Roth
We also point out that minimizing the free energy loss function (22) is equivalent to minimizing Kullback-Leibler (KL) divergence that measures closeness between two distributions. For a statistical physics model the joint probability for spins $\vec{\sigma}=\left\{\sigma_{1},...,\sigma_{N}\right\}$ follows the Boltzmann distribution
@@
IV Low-lying energy states and ground state properties
In this section we present results for the ground state properties and low-lying eigenstates in the energy spectrum for the kagome lattice antiferromagnet clusters with 48 ($4\times 4\times 3$) and 108 ($6\times 6\times 3$) lattice sites and periodic boundary conditions. Whilst the ground state for the smaller lattice size is fully symmetric, we demonstrate further in the section that the ground state for larger cluster with 108 sites is in a nontrivial symmetry sector with nonzero lattice momentum revealing its long-range entanglement,Gioia presence of Dirac points in the energy spectrum and instability towards spinon Cooper pair formation. We note that previous DMRG studies also found indications of the possible spinon Cooper pairing instability. In particular DMRG calculationsYan found that the triplet excitations appear to be composed of two spinons, however could not resolve whether two spinons bind or not on studied cylindrical geometries.
The space group for the $N_{1}\times N_{2}\times 3$ kagome lattice clusters with $N_{1}=N_{2}=L$ and periodic boundary conditions contains $12\times L^{2}$ symmetry operations obtained by combining $L^{2}$ lattice translations and 12 D(6) point group symmetries (rotations and reflections). The space group irreducible representations are usually constructed and described in terms of a set of symmetry related crystal momenta called star, and a subgroup of the point group called little group that leaves these symmetry related crystal momenta invariant.Aroyo
Figure 6: Generators of the kagome lattice point group D(6): (anticlockwise) rotation for $\pi/3$ around the center of the hexagonal plaquette and reflection mapping $y\rightarrow-y$. The table shows group elements of the little group for the crystal momentum $\vec{q}=(0,0)$ ($\Gamma$ high symmetry point in the kagome lattice Brillouin zone) and corresponding irreducible representation characters. D(6) point group contains 12 elements corresponding to 12 point group symmetry operations (6 rotations and 6 reflections). The little group is a subgroup of D(6) that leaves the crystal momentum invariant.
For the cluster with 48 sites ED calculationsLauchli found that the ground state is fully symmetric corresponding to the symmetry sector with all discrete space-group symmetry characters $\chi_{g}=1$. The state is in the total $S=0$ sector with spin parity $P_{\mathbb{Z}_{2}}=1$ As illustrated in FIG. 6, the space-group characters can be obtained from the $\chi_{0}$ characters of the little group for the crytal momentum $\vec{q}=\left(0,0\right)$ of the $\Gamma$ high symmetry point in the kagome lattice Brillouin zone shown in FIG. 4 (further denoted by $\chi_{0}^{0}$).
The first higher energy state in the total $S=0$ sector for 48 sites cluster also has spin parity $P_{\mathbb{Z}_{2}}=1$ and is characterized with the crystal momentum $\vec{q}=(0,0)$ ($\Gamma$ high symmetry point), however with the $\pi$-rotation eigenvalue $-1$ according to ED calculations.Lauchli We find that the corresponding wavefunction transforms as the irreducible representation of the space group constructed from the little group irreducible representations for the crystal momentum $\vec{k}=\left(0,0\right)$ described by characters $\chi_{2}$ in FIG. 6. The space-group characters in this symmetry sector are further denoted by $\chi_{0}^{1}$.
Furthermore, ED studies found that the lowest energy state in the total $S=1$ sector for 48 sites cluster has spin parity $P_{\mathbb{Z}_{2}}=-1$, $\pi$-rotation eigenvalue $-1$ and finite crystal momentum $\vec{q}=\left(\pi,\pi/\sqrt{3}\right)$.Lauchli The wavefunction is found to transform according to the irreducible representation of the space group constructed from the little group irreducible representation for the crystal momentum $\vec{q}=\left(\pi,\pi/\sqrt{3}\right)$ ($M$ high symmetry point in the kagome lattice Brillouin zone shown in FIG. 4) described by characters $\chi_{2}$ in FIG. 7. We denote space-group characters for this symmetry sector by $\chi_{1}^{0}$.
To check accuracy of the GCNN and VMC approach we first calculate the ground state energy and singlet and triplet energy gaps for 48 sites kagome lattice cluster. The singlet excitation gap $\Delta_{s}=E_{1}(S=0)-E_{0}(S=0)$ is defined as energy difference of the two lowest energy states in the total spin $S=0$ sector, whilst the triplet spin excitation gap $\Delta_{t}=E_{0}(S=1)-E_{0}(S=0)$ can be obtained as energy difference of the lowest energy states in the total spin $S=1$ and $S=0$ sectors.
Figure 7: Group elements of the little group for the crystal momentum $\vec{q}=(\pi,\pi/\sqrt{3})$ (M high symmetry point in the kagome lattice Brillouin zone) and corresponding irreducible representation characters. Here $R$ is (anticlockwise) $\pi/3$ rotation around the center of the hexagonal plaquette and $\tau$ is reflection about the reflection line with angle $-\pi/3$ with respect to the $x$-axis.
First the ground state energy is found by minimizing the free energy loss (22) with fully symmetric wave-function ansatz with space-group characters $\chi_{0}^{0}$ and spin parity $P=1$ using SR optimization algorithm and the low-lying excited states are further found using transfer learning approach. Namely to stabilize and expedite training within SR optimization for the lowest energy states in the symmetry sectors described by space-group characters $\chi_{0}^{1}$ and $\chi_{1}^{1}$, GCNN ansätze for those states are taken to have the same architecture as for the ground state and the optimized ground state parameters are used as initial guess for the parameters of the ansätze. The ansätze therefore have the same number of layers and the same number of features in each layer as the ground state ansatz, however with different space group irreducible representation characters that reflect symmetry properties of a particular low-lying energy state. The ansätze can in general also have different spin parity $P_{\mathbb{Z}_{2}}$ equal to $1$ (for $\chi_{0}^{1}$ symmetry sector) or $-1$ (for $\chi_{1}^{0}$ symmetry sector). Within this approach initial guess for a particular low-lying energy state wave function already has low variational energy and typical correlations characteristic for the low-energy eigenstates resulting in more stable training yielding more accurate results for the energy eigenvalues.

### v2

@@ Appendix A Group convolutional neural networks
and $\vec{\sigma}=\left\{\sigma(\vec{r}_{1}),...,\sigma(\vec{r}_{N})\right\}$ with $\vec{r}_{1},...,\vec{r}_{N}$ being positions of the lattice sites. Spin configurations $|\vec{\sigma}\rangle$ are the network inputs, and the complex coefficients $\psi(\vec{\sigma})$ are the network outputs as illustrated in FIG. 11.
@@
All $N_{l}$ hidden layers are followed by the $\Gamma$ activation function. To obtain $\psi(\sigma)$ in Eq. (27) the final output layer exponentiates complex-valued feature maps composed of two real-valued feature maps, calculates sum of these exponentiated values, and projects the sum on a particular irreducible representation of the symmetry group $G$ that corresponds to the symmetry properties of the ansatz wavefuntion:
@@
Within group theory representations are eigenfunctions of the symmetry group operations. Irreducible representations are eigenfunctions of all group operations whilst reducible representations are eigenfunctions only of a subset of these operations. Each irreducible representation has different set of eigenvalues. These eigenvalues are characters $\chi_{g}$ in the Eq. (34). For non-degenerate symmetry groups where every irreducible representation consists of one eigenfunction, characters are $+1$ or $-1$, whilst for degenerate symmetry groups characters can differ from ±$\pm 1$.
@@
It should be noted that the described procedure of constructing the wave-function ansätze equivariant with respect to a discrete symmetry group cannot straightforwardly be generalized to continuous symmetry groups like $SU(2)$ spin rotation symmetry. Finding efficient ways to introduce continuous symmetries is still an open research problem with one of the possible solutions being sampling in the basis of irreducible representations instead of spins.Vieijra ; Vieijra2 ; Misawa
Appendix B Training scheme and wave-function optimization
The goal of VMC is to find the optimal set of (complex) parameters $\vec{\alpha}=\left\{\alpha_{1},...,\alpha_{N_{p}}\right\}$ in an ansatz wavefunction of the form (27) that minimizes a suitably chosen loss function. The standard optimization technique is the gradient descent (GD) method within which the parameters are updated according to the gradients of the chosen loss function with the iterative update rule
@@ Appendix B Training scheme and wave-function optimization
where $t$ denotes iteration, $\eta$ is the learning rate (the step size) and $\mathcal{L}$ is a loss function. All numerical methods that attempt to find a global minimum using local information for iterative parameter updates generally exhibit problems related to the local minima in the loss function landscape and slow convergence. Various stochastic methods are therefore introduced to help mentioned numerical algorithms to escape local minima and improve convergence, for example AdamKingma and natural gradient descent (NGD) method.Amari ; Park ; Martens ; Dong ; Park2 ; Stokes2
In the Eq. (38) -$\vec{\nabla}_{\vec{\alpha}}\mathcal{L}$ represent directions in which the parameters can be updated to obtain the largest decrease of the loss function. The parameters are adjusted by a small step length, i.e. moved for a small distance in the parameter space, in the direction determined by the gradient to optimize the loss function. The standard GD and NGD methods differ by how a small distance is defined. Whilst in the standard GD distance is Euclidean distance that corresponds to the flat metric on the parameter manifold, NGD replaces the flat metric by a generally non-flat metric. NGD method takes into account that changes of all parameters can not be treated equally and that it is not always correct to define distance in terms of how much the parameters are adjusted. Instead a change for each parameter has to be scaled according to the effect of the change on the neural network output distribution.
@@
where $\Delta\tau$ is a small imaginary time step that corresponds to the learning rate $\eta$ in the Eq. (39). After each imaginary time step $\Delta\tau$ the wavefunction $|\psi^{\prime}\left(\vec{\alpha}\left(\tau\right)\right)\rangle$ is projected to the variational subspace by minimizing $d_{FS}^{2}$ with $|\phi\rangle=|\psi\left(\vec{\alpha}\left(\tau\right)\right)\rangle+\Delta\tau\partial_{\tau}\psi\left(\vec{\alpha}\left(\tau\right)\right)$ and $|\phi^{\prime}\rangle=|\psi^{\prime}\left(\vec{\alpha}\left(\tau\right)\right)\rangle$ in the Eq. (40) to obtain a new set of parameters $\vec{\alpha}_{t+1}\equiv\vec{\alpha}\left(\tau+\Delta\tau\right)$:
@@
The loss function in VMC can for example be the variational energy:
@@
where $H$ is the Hamiltonian of the system or free energy at temperature $T$:
@@
where the entropy is defined asRoth ; Roth2 ; Roth3 ; Wu ; Hibat-Allah2 ; Hibat-Allah3
@@
and $\psi\left(\vec{\sigma},\vec{\alpha}\right)$ is an unnormalized neural-network ansatz wavefunction.
Minimizing the free energy loss function (45) corresponds to minimizing Kullback-Leibler (KL) divergence that measures closeness between two distributions. For a statistical physics model the joint probability for spins $\vec{\sigma}=\left\{\sigma_{1},...,\sigma_{N}\right\}$ follows the Boltzmann distribution